OpenStudy (simplymarie_x):

Explain how to find the inverse of f(x) = (x-1)/4

OpenStudy (harman.singh):

You can generally solve for inverse of a function using the following steps: 1) Replace f(x) with y 2) Switch x's and y's 3) Solve for y 4) Replace 'y' with 'f^-1(x)'

OpenStudy (518nad):

y=x*4 +1 y=4x+1

OpenStudy (518nad):

you see from PEMDAS which tells your the order of operations to perform, has to be undone to get what the input was for a given output.

OpenStudy (simplymarie_x):

Is this right: ? if g(x) is the inverse of f(x), then, by definition, f(g(x)) = x f(x) = (x-1)/4 f(g(x)) = (g(x)-1)/4 = x g(x) = 4x + 1 But, since g(x) is the inverse of f(x), f^-1(x) = g(x) = 4x+1

OpenStudy (518nad):

so if you look at your expression f(x)=(x-1)/4 the first thing you did was -1, then we /4 so to undo this we look at the last operation /4, that becomes *4 then -1 becomes +1 f-1(x) = 4*x + 1

OpenStudy (518nad):

yeah you did it right

OpenStudy (simplymarie_x):

Thank you!

OpenStudy (518nad):

I like how you use f(g(x)) it's very clean